### 28 Feb 2024 v1

# Unitary coupled-cluster singles and doubles (UCCSD) calculations in conjunction with fragment molecular orbital (FMO) scheme Kenji Sugisaki , Tatsuya Nakano, Yuji Mochizuki

Abstract
The fragment molecular orbital (FMO) method is one of the popular methods to efficiently treat macromolecular systems by dividing the system of interest into small fragments based on embedding under the electrostatic potential (ESP). Such a fragmentation method has the potential advantage of making the circuit flat in quantum chemical calculations on quantum computers. In this study, we used a GPU-accelerated quantum simulator (cuQuantum) to perform the electron correlation part of the FMO calculation as a unitary coupled-cluster singles and doubles (UCCSD) with the variational quantum eigensolver (VQE), using hydrogen-bonded systems (FH) and (FH)-HO as testbeds. From the numerical simulations, we found that the Trotter decomposition of the UCCSD ansatz breaks the orbital-invariance condition, and it can yield different total energies for symmetrically equivalent molecules. We also observed that the Trotterized UCCSD does not satisfy the size-consistency, which is an essential requirement in the FMO scheme, when the molecular orbitals delocalized to dimers are used. The GPU acceleration was substantial for the simulations with larger numbers of qubits, and it was about a factor of 5.8–7.5 for 18 qubit systems.
Keywords
Fragment Molecular Orbital (FMO), Unitary Coupled-Cluster (UCC), cuQuantum, GPU, Size-Consistency, Orbital-Invariance, Trotter Error
1. Introduction
Starting with the seminal work of Aspuru-Guzik et al. [1], quantum chemical computation has been actively explored and developed as a promising application area for quantum computers [2, 3, 4, 5, 6], where the potential applicability to huge-scale configuration interactions [7] is attractive. In practice, however, the development of computational methods and algorithms using quantum simulators is currently more mainstream than the use of actual devices. For noisy intermediate-scale quantum (NISQ) computers, the unitary coupled-cluster singles and doubles (UCCSD) [8, 9] has been used for relatively small molecules in conjunction with the variational quantum eigensolver (VQE) [10, 11, 12, 13], and this VQE-UCCSD scheme has been extended to multi-reference cases, e.g., Refs. [14, 15, 16, 17]. In addition, GPU-accelerated simulators such as cuQuantum [18] have attracted considerable interest due to its pronounced performance [19].
In another direction, the so-called problem decomposition approach has been introduced to shallow the circuit depth [20] while avoiding the effects of noise. Note that such an approach is rather common for large molecules (like proteins) as the fragmentation-based methods [21, 22, 23]. The introduction of problem decomposition to quantum computation was pioneered by Yamazaki et al [24], who compared three methods of fragment molecular orbital (FMO) [25], divide-and-conquer (DC) [26], and density matrix embedding theory (DMET) [27]. Recently, the FMO calculations with VQE-UCCSD for hydrogen clusters have been reported [28].
In this study, we applied the VQE-UCCSD scheme [10, 11, 12, 13] to compute the electron correlation part of FMO calculations [25, 29, 30] for a couple of hydrogen-bonded systems, (FH) and (FH)-HO. For execution, the cuQuantum simulator [18] was used as done in the previous study [19]. In the following we describe the calculation method, the results and the discussion.
2. Method of calculation
2.1. FMO scheme and program
The scheme of the basic two-body FMO calculation [25, 29, 30] is summarized as follows. The first step is to determine the molecular orbital and electron density of each monomer by the Hartree–Fock (HF) approximation [31] under a given basis function, while self-consistently imposing an electrostatic potential (ESP) on each other. The set of ESPs of the monomers is to be determined until the monomer self-consistent-charge (SCC) condition is satisfied by iterations. This allows the polarization of each monomer to be taken into account. In the next step, the monomer-determined ESP is used to calculate the HF for the dimer; no SCC condition is imposed. The dimer calculation takes into account the delocalization of electrons between the monomers. From the sum of the HF energies of the monomer and the dimer, the two-body FMO energy of the system of interest is given as in Eq. (1)
Indices of and specify the respective monomers, and is the number of fragments.
Electron correlation calculations, such as second-order Møller–Plesset perturbations (MP2) [31], are performed after the HF calculations for each monomer are complete and after the individual HF calculations for each dimer are complete. The correlation energy correction is done in an additive manner as in Eq. (1). The introduction of electron correlations is essential to improve quantitatively by incorporating dispersion stabilization and reducing excess ionicities of the HF description. As described in Ref. [31], both size-consistency and orbital-invariance are crucial requirements in the correlated methods. This is obviously true for the FMO scheme based on Eq. (1).
Currently, GAMESS-US [32], PAICS [33], and ABINIT-MP [34, 35] are the available programs that can perform FMO calculations including electron correlation correction by MP2. Besides the MP2 capability [36, 37, 38], ABINIT-MP is unique in supporting higher correlated calculations [31, 39] on-the-fly; from the third-order MP (MP3) [40] to coupled-cluster singles and doubles including perturbative triples (CCSD(T)) [41] are supported.
2.2. Preparation of molecular integrals under FMO scheme
The geometries of (FH) (under C symmetry) and (FH)-HO (C symmetry) were optimized by the GAUSSIAN16W program [42] at the level of B3LYP [43] corrected with the empirical dispersion [44] with the 6-31+G(d’,p’) basis set [45]. The resulting Cartesian coordinates are listed in Table 1 and illustrated in Figure 1.
| Seq. | Frag. | Elem. | x | y | z |
| 1 | 1 | F | 0.779023 | 0.287467 | 0.000000 |
| 2 | 1 | H | 0.000000 | 0.807300 | 0.000000 |
| 3 | 2 | F | 1.361653 | 1.874668 | 0.000000 |
| 4 | 2 | H | 1.330503 | 2.801407 | 0.000000 |
| 5 | 3 | F | 0.648089 | 2.399606 | 0.000000 |
| 6 | 3 | H | 0.741364 | 1.471463 | 0.000000 |
| (FH)-HO | |||||
| 1 | 1 | O | 0.000000 | 0.000000 | 1.200039 |
| 2 | 1 | H | 0.000000 | 0.771251 | 1.779594 |
| 3 | 1 | H | 0.000000 | 0.771251 | 1.779594 |
| 4 | 2 | F | 0.000000 | 2.034006 | 0.694256 |
| 5 | 2 | H | 0.000000 | 1.179139 | 0.331451 |
| 6 | 3 | F | 0.000000 | 2.034006 | 0.694256 |
| 7 | 3 | H | 0.000000 | 1.179139 | 0.331451 |
For (FH) and (FH)-HO, the FMO calculations were performed with the STO-3G minimal basis set [46], where we used a local version of ABINIT-MP, which dumped the integral list of basis functions and the converged MO coefficients (at the FMO-HF level) of monomers and dimers as separate files. These data were transformed by a small Fortran program into molecular integrals for the second-quantized Hamiltonian used to run VQE-UCCSD, expressed as
Indices of and cover the correlating orbital space, and and are the transformed one- and two-electron integrals. Note that there is some degree of locality of monomer orbitals in dimer orbitals with respect to the occupied space for (FH) and also that there is the symmetric delocalization for the FH dimer in (FH)-HO. Due to a proof-of-concept (PoC) phase of this study, the FMO calculations by ABINIT-MP were done in a separate step from the quantum calculations described in the next subsection. The 1s orbitals of fluorine and oxygen were frozen [47] for in Eq. (2), which is a good approximation to save on the number of qubits [48, 49]. The number of correlated electrons for dimers was thus 16. For comparison with the VQE-UCCSD correlation energies, the usual FMO-MP2 and FMO-CCSD(T) calculations were also performed by ABINIT-MP. These calculations were completed in less than 1 second on a single core of Intel Xeon processor.
2.3. Set-up of VQE-UCCSD calculation
VQE is a quantum–classical hybrid algorithm and it has been proposed to solve quantum chemistry problems using NISQ devices [10, 11]. In VQE, an approximate wave function is generated by using a parameterized quantum circuit (PQC) defined by an “ansatz”, and the expectation value of the qubit Hamiltonian obtained by applying the fermion–qubit transformation to the second-quantized Hamiltonian given in Eq. (2) is computed statistically, by repeatedly executing the quantum circuit and collecting the measurement results. The classical computer then execute a variational optimization of the parameters in PQC. These steps are iterated until convergence.
Various types of ansatzes have been proposed and studied for quantum chemical calculations [50]. Here, we adopted the UCCSD ansatz defined in Eqs. (3) and (4), because it is a chemically motivated ansatz and it can give very accurate correlation energies.
Here, we used the indices and for the occupied spin orbitals and and for the unoccupied orbitals of the HF wave function . To accelerate the VQE simulations, we adopted the following techniques: (1) Using the symmetry conserving Bravyi–Kitaev transformation (SCBKT) [51] to reduce two qubits in the simulation, (2) using the MP3 and the MP2 excitation amplitudes as the initial guess of the and , respectively [17], and (3) GPU-based numerical simulations. The number of qubits for the VQE-UCCSD simulations was 8 and 18 for monomers and dimers, respectively, in (FH), and 10 for monomer “1” and 20 for dimers “21” and “31” in (FH)-HO. The VQE-UCCSD simulation program was developed by us, by using Python3 with OpenFermion [52], Cirq [53], and cuQuantum [18] libraries.
In the implementation of the UCCSD quantum circuit, we adopted the first-order Trotter decomposition with one Trotter slice in conjunction with the magnitude ordering [54] of the cluster operators. For the variational optimization of the excitation amplitudes, we examined COBYLA [55] and Powell [56] algorithms; the corresponding labels are shortly denoted as UCCSD:CB and UCCSD:PW, respectively (see Tables 2 and 3). For comparison, the calculation of the complete active space configuration interaction (CAS-CI) was performed to obtain the exact correlation energy in the orbital space of STO-3G. The numerical simulations for (FH) and (FH)-HO were carried out on the Supercomputer ‘Flow’ Type-II subsystem at Nagoya University and on the in-house NVIDIA DGX H100 system, respectively.
3. Results and discussion
3.1. Energies and timings
The correlation energies for (FH) are summarized in Table 2. Compared to MP2, CCSD has a significantly lower energy, and CCSD(T) gives values close to CAS-CI, as expected. The quantum calculation for UCCSD is obtained by two different solvers, COBYLA and Powell. Powell gave lower energies than COBYLA, but the number of function evaluations (total energy calculations) required in Powell is about 1.6–2.8 times greater than in COBYLA (115, 130, 119, 1762, 1963, and 1910 for “1”, “2”, “3”, “21”, “31”, and “32”, respectively, for COBYLA, and 225, 257, 193, 4783, 4765, and 5408 for Powell).
| Unit | MP2 | CCSD | CCSD(T) | UCCSD:CB | UCCSD:PW | CAS-CI |
| Monomer | ||||||
| “1” | 0.017933 | 0.026945 | 0.026945 | 0.026884 | 0.026914 | 0.026945 |
| “2” | 0.017526 | 0.026216 | 0.026216 | 0.026164 | 0.026192 | 0.026216 |
| “3” | 0.017933 | 0.026929 | 0.026929 | 0.026899 | 0.026875 | 0.026929 |
| Dimer | ||||||
| “21” | 0.035493 | 0.051856 | 0.051933 | 0.049880 | 0.050452 | 0.051963 |
| “31” | 0.035980 | 0.052778 | 0.052850 | 0.051554 | 0.051527 | 0.052879 |
| “32” | 0.035446 | 0.053122 | 0.053123 | 0.051952 | 0.053067 | 0.053124 |
| Sum. | 0.053527 | 0.077666 | 0.077816 | 0.073439 | 0.075065 | 0.077876 |
Table 3 summarizes the results for the (FH)-HO correlation energies. A checkpoint here is whether the equivalence symmetries (monomers “2” and “3” / dimers “21” and “31”) are satisfied, and the usual MP2, CCSD, CCSD(T), and CAS-CI results all satisfy this requirement. The trend of the correlation energies by these methods is the same as in (FH). On the other hand, the UCCSD results (of both COBYLA and Powell) unfortunately do not satisfy symmetry, as the difference is seen to five decimal places for monomers and three decimal places for dimers. From a chemical precision point of view, it seems problematic that the effect is seen to three decimal places. Furthermore, this issue of broken equivalence should be kept in mind not only for FMO, but for all approaches of fragmentation-oriented methods [21, 22, 23]. This problem of VQE-UCCSD is related to the Trotter error and is discussed in the next section.
| Unit | MP2 | CCSD | CCSD(T) | UCCSD:CB | UCCSD:PW | CAS-CI |
| Monomer | ||||||
| “1” | 0.035370 | 0.049321 | 0.049394 | 0.049242 | 0.049231 | 0.049445 |
| “2” | 0.017810 | 0.026705 | 0.026705 | 0.026608 | 0.026692 | 0.026705 |
| “3” | 0.017810 | 0.026705 | 0.026705 | 0.026677 | 0.026659 | 0.026705 |
| Dimer | ||||||
| “21” | 0.053486 | 0.075392 | 0.075526 | 0.069313 | 0.071421 | 0.075600 |
| “31” | 0.053486 | 0.075392 | 0.075526 | 0.071847 | 0.072544 | 0.075600 |
| “32” | 0.035790 | 0.053549 | 0.053553 | 0.053230 | 0.053207 | 0.053549 |
| Sum. | 0.071770 | 0.101602 | 0.101801 | 0.091863 | 0.094590 | 0.101895 |
The cuQuantum quantum simulator was used in this VQE-UCCSD computation. Table 4 summarizes the timings of the UCCSD jobs of (FH) on the ‘Flow’ Type-II subsystem with and without GPU. The GPU acceleration was about a factor of 1.8–2.5 and 5.8–7.5 for monomers and dimers, respectively. GPU acceleration is substantial for larger systems, but the speedup is less significant compared with our previous study [19]. This is because the VQE job needs a lot of time for pre-processing and post-processing, and these parts cannot be accelerated by cuQuantum. Considering that a normal FMO-CCSD(T)/STO-3G calculation takes less than 1 second to complete, there is a speed difference of the order of the fourth power of 10 if the correlation part is due to VQE-UCCSD at this time. Note that the present VQE-UCCSD was run on a classical computer, where computational time grows exponentially with the number of qubits. Anyway, as in the previous report [19], GPU acceleration with cuQuantum is essential for quantum simulations.
| UCCSD:CB | UCCSD:PW | |||||
| Unit | with GPU | without GPU | Acceleration | with GPU | without GPU | Acceleration |
| Monomer | ||||||
| “1” | 14.1 | 30.1 | 2.13 | 25.4 | 52.3 | 2.06 |
| “2” | 16.1 | 29.1 | 1.80 | 27.9 | 68.9 | 2.47 |
| “3” | 14.8 | 31.5 | 2.13 | 22.2 | 48.4 | 2.18 |
| Dimer | ||||||
| “21” | 11901.3 | 70531.8 | 5.93 | 31276.4 | 216982.7 | 6.94 |
| “31” | 12871.9 | 74618.2 | 5.80 | 30315.4 | 210808.9 | 6.95 |
| “32” | 11209.1 | 83915.6 | 7.49 | 27740.4 | 203298.3 | 7.33 |
3.2. Relation with Trotter error
It is interesting to note that the monomers “2” and “3” of (FH)-HO are symmetrically equivalent, but VQE-UCCSD yields different correlation energies. The HF molecular orbitals of the monomers “2” and “3” are illustrated in Figure 2. We found that relative phase of the second and the fourth molecular orbitals is different between monomers “2” and “3”, which causes differences in the absolute sign of excitation amplitudes and . As a result, the quantum state corresponding to the UCCSD wave function is not identical for monomers “2” and “3”, and the Trotter error appears in a different way. This fact is confirmed by performing the UCCSD calculations without Trotterization, using the expm_multiply function in the SciPy library [57], which allows us to compute the action of the matrix exponential of on . In this case, the calculated correlation energies of the monomers “2” and “3” are exactly the same: Hartree.
Since the Trotter errors appears in an unexpected way, we further investigated about the relationship between Trotter errors and size-consistency, which is an essential condition in the FMO framework as mentioned earlier. Here we focused on the tetrahydrogen (4H) cluster [58] in a square coordinate with R(H–H) = 1.0583 Å (2.0 Bohr) as the monomer, because the Trotter error becomes more significant when the HF is not a good approximation of the ground-state wave function and the UCCSD wave function has large excitation amplitudes. This system is also suitable because the HF molecular orbitals are completely defined by point-group symmetry. In the dimer (8H) calculations, two 4H clusters were placed to form a cuboid, with the inter-monomer distance being 100 Å. Two types of molecular orbitals are examined in the dimer calculations: Completely delocalized HF canonical orbitals in D point group (denoted as CMO) and orbitals localized on the monomers (denoted as LMO). In the total energy calculations using UCCSD, we used cuQuantum-based quantum circuit simulations with Trotter decomposition and Trotter-free calculations using the expm_multiply function in SciPy. The results are summarized in Table 5.
| System | Trotter decomposition | |
| Monomer | No | 0.8118 |
| Dimer (LMO) | No | 1.6236 |
| Dimer (CMO) | No | 1.6234 |
| Monomer | Yes | 0.8102 |
| Dimer (LMO) | Yes | 1.6207 |
| Dimer (CMO) | Yes | 5.0319 |
From Table 5, the UCCSD calculations without Trotter decomposition yield almost the same values for both LMO and CMO-based calculations, and the values of the dimer are twice of that of the monomer; i.e., Trotter-free UCCSD satisfies the size-consistency condition. Small differences in the values of dimer with CMO and LMO are due to rounding errors in the AO MO transformation. In contrast, when the Trotter decomposition is used to construct the UCCSD quantum circuit, the of dimer with CMO is significantly larger than the . Since the value for dimer with LMO is approximately double of the of monomer, we concluded that size-consistency condition of the VQE-UCCSD can be maintained when the molecular orbitals localized on each monomer are used in the calculations. This observation is very important for the application of the VQE-UCCSD method to large systems. In the present FMO-UCCSD calculations the molecular orbitals used for the dimers are not completely localized on the monomers, and the size-consistency condition may be broken to some extent. We expect that the use of localized molecular orbitals will improve the correlation energy of the dimers and the sum of the correlation energies, and the FMO-UCCSD calculations with the localized orbitals are in progress. It should be noted that the dependence of the Trotter error on the locality of the molecular orbitals was investigated by Babbush and coworkers [59], and they reported that the Trotter error is larger for localized orbital basis than for canonical orbitals and natural orbitals. The reported study focused only on single molecule (monomer), and our discussions are based on comparing the energies of monomers and a dimer. Our results do not contradict the previous study.
4. Summary
We have performed VQE-UCCSD calculations [9, 10, 11, 12, 13] using the cuQuantum simulator [18] in conjunction with the FMO calculations for the (FH) and (FH)-HO systems. The STO-3G minimal basis set [46] was used and the frozen-core restriction [47] was imposed. By combining with symmetry conserving Bravyi–Kitaev transformation [51], we can simulate the HFHO dimer with 20 qubits. Both COBYLA and Powell methods were used for VQE optimization, with the latter usually providing better energies but requires more function calls. (FH)-HO system has two symmetrically equivalent FH molecules that should have the same energies, but the Trotterized UCCSD does not satisfy symmetry and yields different energies. The difference in correlation energies between the dimers “21” and “31” is on the order of 1 kcal mol. Size-consistency of the Trotterized UCCSD is also studied numerically using 4H and 8H model clusters. We found that the size-consistency condition can be broken when the molecular orbitals delocalized to the dimer are used for the calculation, and using molecular orbitals localized to the monomers is essential to satisfy the size-consistency. These findings on the relationship between size-consistency and orbital-invariance [31, 39] and the error in the Trotter decomposition are very important not only for FMO-based quantum chemical calculations and other fragmentation-oriented methods [21, 22, 23] but also for VQE-UCCSD in general. Possible future countermeasures include choosing an appropriate optimization algorithm, setting a stricter convergence threshold, using finer Trotter slices, or adopting higher-order Trotter decomposition.
The GPU acceleration was found to be 7.49 and 7.33 with COBYLA and Powell algorithms, respectively, for the dimer “32” of (FH). The usefulness of GPUs has attracted much attention in various fields of quantum computation [60], and quantum chemistry is an example where the acceleration effect is significant [61], including in this case; even with GPU acceleration, it still takes orders of magnitude longer than a regular FMO-CCSD(T) calculation [41]. Recently, an example of large-scale quantum computation with adamantanes has been reported using VQE [62]. Following these trends, we will perform larger FMO-UCCSD computations on upcoming GPU environments.
5. Acknowledgement
All VQE-UCCSD computations with cuQuantum on the ‘Flow’ Type-II subsystem at the Information Technology Center of Nagoya University were performed under the JHPCN Joint Research Projects (jh230001 subject by YM). KS and YM would like to thank Profs. Takahiro Katagiri (Nagoya University) and Satoshi Ohshima (Kyushu University) for their encouragement. YM was also supported by Rikkyo SFR. KS acknowledges the support from Quantum Leap Flagship Program (Grant No. JPMXS0120319794) from MEXT, Japan, Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from JST, Japan, and KAKENHI Transformative Research Area B (23H03819) and Scientific Research C (21K03407) from JSPS, Japan.